Exercise 3. (22 pts.) (Rabbits vs. Sheep) a)For each of the following ”rabbits vs. sheep” problems with x, y ≥ 0. Find the fixed points, classify them (type and stability), and find the eigenvalues and eigenvectors. Then sketch (by hand) a plausible phase portrait indicating nullclines, all relevant trajectories, and indicate all the different basins of attraction. Finally interpret the behavior of the population of the rabbits and sheep for the different systems (compare the three systems)

• x˙ = x(3 − 2x − 2y), ˙y = y(2 − x − y)
Answer #1

Given system. iu (3-2x-2y) e 30-24²-2my. = 9(2-*-y). 2y-ay-y for the fixed pts. x = 0 = ý. ..(3-2x-2y) > 0. y(2-x-y). When 12Now, consider the fixed pt. (0,2). A2 / - 01 2 eigenvalues are at ,-2. 1-2 -2%. (0, 2) is a stable saadados a node eigenvertacedex 5 al x+y=2. x+y = 3 n. 20 Y=0 The nullclinex are x+y a 2/3, n+y=2. Phase portrait Sheep Yabbit- The chase portrait inteGriven system. ni = x(3-28-4) y = y(2-x-4). for the fixed pts. ģ= 0 = ý. 2(3-2x-4) = 0 = 9 (2-3-4). When n2o. & 7 y(2-x-4) 20Now, consider the fixed pt. (0,2). A o -2 1-2 eigenvalues 10.2) is a . are 1-2 unstable saddle point eigenvertar are (-2), (iPhase portrait So any open neighbourhood of (1,1) will be the basin of attraction i.e. the positive quadraint of the IRL is t